OpenStudy (itsmichelle29):

How can1/3 x − 2 = 1/4 x + 11 be set up as a system of equations?

OpenStudy (itsmichelle29):

help i medal and fan

Vocaloid (vocaloid):

use a second variable, y y = 1/3x - 2 y = 1/4x + 11

OpenStudy (itsmichelle29):

3y − x = −6 4y − x = 44 3y + x = −6 4y + x = 44 3y − 3x = −6 4y − 4x = 44 3y + 3x = −6 4y + 4x = 44

OpenStudy (itsmichelle29):

those are the answers

OpenStudy (itsmichelle29):

@Vocaloid

Vocaloid (vocaloid):

well, we can start off with y = 1/3x - 2 y = 1/4x + 11 we can multiply the first equation by 3, and the second equation by 4, giving us 3y = x - 6 4y = x + 44 we can then subtract x from both sides, giving us 3y − x = −6 4y − x = 44

OpenStudy (itsmichelle29):

Thanks :)